OpenStudy (anonymous):

Which of the following is the smallest volume? 500 mL 2500 cm3 5.5 x 10-1 L 25 m3

OpenStudy (anonymous):

You'll need your SI prefix conversion ability, plus the conversion between cm^3 and mL (1 cm^3 = 1 mL) to put them all in the same unit. Let's pick mL. 1) 500 mL = 500 mL. That was easy! 2) 2500 cm^3 = 2500 mL (using 1 cm^3 = 1 mL). 3) 5.5 x 10^-1 L = 0.55 L = 0.55 L * (1000 mL/L) = 550 mL 4) 25 m^3 = 25 m^3 (100 cm/m)^3 = 25 x 10^6 cm^3 = 25 x 10^6 mL. Now you can choose pretty easily.

OpenStudy (anonymous):

@mangorox so it would be 500 ml is the smallest?

OpenStudy (anonymous):

Yes, it would be 500 ml as the smallest.